Area bounded by the curve y = sinx between x = 0 and x = 2p is
(A) 2sq. units
(B) 4 sq. units
(C) 8sq. units
(D) 16sq. units

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Area ‘A’:
A = \int_{0}^{\pi }sinx.dx + |\int_{\pi }^{2\pi }sinx.dx|
A = (-cosx)_{0}^{\pi } + |(-cosx)_{\pi }^{2\pi }| = 2 + |-2| = 4
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty